OpenStudy (anonymous):

Simplify 2 to the 5th over 3 squared all raised to the 4th power. 2 to the 20th over 3 to the 8th 2 to the 9th over 3 to the 8th 8 to the 5th over 12 to the 2nd 2 over 3 to the 2nd

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ \left(\frac{2^5}{3^2} \right)^4\]

OpenStudy (anonymous):

thatscorrect

OpenStudy (phi):

so you have "something" multiplied by itself 4 times

OpenStudy (anonymous):

ok

OpenStudy (phi):

what is multiplied by itself 4 times?

OpenStudy (anonymous):

the 5?

OpenStudy (anonymous):

or 2

OpenStudy (phi):

everything in the parens is treated as one thing, that is multiplied by itself 4 times.

OpenStudy (anonymous):

so the 2 5 3 and 2 are multiplied by 4?

OpenStudy (phi):

you make a 4 copies of \(\left(\frac{2^5}{3^2} \right) \) and multiply them together.

OpenStudy (anonymous):

woah

OpenStudy (anonymous):

thats alot

OpenStudy (phi):

do you know how to multiply fractions? top times top and bottom times bottom

OpenStudy (anonymous):

a little bit im pretty bad when it comes to fractions

OpenStudy (phi):

care to figure out the top ? (it's 2^5 times itself 4 times)

OpenStudy (anonymous):

mmmmh

OpenStudy (anonymous):

idk

OpenStudy (phi):

do you know what 2^5 means ? 2*2*2*2*2 right ?

OpenStudy (anonymous):

ohh woops

OpenStudy (anonymous):

ya so we do that

OpenStudy (anonymous):

4 times?

OpenStudy (phi):

yes

OpenStudy (anonymous):

ok 1 sec

OpenStudy (phi):

just "glue them together" and count the 2's

OpenStudy (anonymous):

2x2x2x2x2x2x2x2x2x2x2x2x2x2x2x2x2x2x2x2 20? xD

OpenStudy (phi):

yes 2^20 now the bottom

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x3x3x3

OpenStudy (anonymous):

right?

OpenStudy (phi):

the bottom is 3^2 times itself 4 times

OpenStudy (anonymous):

oh i need 2 more sorry

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

isnt that correct or would it be 3x2x3x2x3x2x3x2

OpenStudy (phi):

just 3 times itself

OpenStudy (anonymous):

3x3x3x3x3x3

OpenStudy (phi):

3^2 is 3*3 we multiply 3*3 by itself 4 times yes 3^8

OpenStudy (phi):

so the answer is 2^20/ 3^8

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

thats wasnt that bad

OpenStudy (phi):

practice helps

OpenStudy (anonymous):

yup ty very much

OpenStudy (anonymous):

ok 1 more and we will be done :DD

OpenStudy (phi):

ok, I'll get there. but try to do some of the work.